# Distributed Model Checking in Graphs Classes

### v1

@@ 3.1.2 Quantifier Elimination of lca-Reduced Formulas
$\beta(\overline{x})=\bigvee_{q\in[p]}\bigg{[}\textsf{lab}_{q}(x_{1}))\wedge\bigvee_{S\in K(p,q-1)}\bigg{(}\bigwedge_{i\in S}\bigvee_{j\in X_{i}}\textsf{lab}_{\textsf{good}}(x_{j})\wedge\bigwedge_{i\notin S}\bigwedge_{j\in X_{i}}\neg\textsf{lab}_{\textsf{good}}(x_{j})\bigg{)}\bigg{]}.$
Case 3: $\delta(1,k+1)=-1$.
In this case, the node $y$ in $F_{\psi}$ has no common ancestors with any of the nodes $x_{1},\dots,x_{k}$. In other words, $y$ is in a tree of $F_{\psi}$ that is different from any of the trees containing $x_{1},\dots,x_{k}$. Let $(F,L)\in\mathcal{F}_{d}[\Lambda]$. Like in previous cases, a node $u$ at depth $h_{y}$ is called a candidate if $L(u)=\gamma(k+1)$. A root $r$ of $F$ is said to be active if some descendant of $r$ is a candidate node. A node is called good if it is a descendant of an active root.
@@
Let $\overline{v}=(v_{1},\dots,v_{k})$ be a set of vertices of $F$, and let $\rho(F)$ be the number of active roots in $F$. Similarly, let $\rho(F,\overline{v})$ be the number of active roots in $F[\overline{v}]$. It follows that $(F,L)\models\exists y,\psi(\overline{v},y)$ if and only if $F[\overline{v}]$ is isomorphic to $F_{\xi}$, $L(v_{i})=\gamma(i)$ for each $i\in[k+1]$, and $F$ contains an active root not contained in $F[\overline{v}]$. The first two conditions can be checked using $\xi$, and the latter is equivalent to $\rho(F)>\rho(F,\overline{v})$.
@@
Let $\widehat{\Lambda}=\Lambda\cup[0,k]\cup\{\infty,\textsf{good}\}$ be a set of labels, and let us consider the $\widehat{\Lambda}^{\psi}$-labeling $\widehat{L}^{\psi}$ of $F$ defined as follows. Initially, for every node $u$ in $F$, let us set $\widehat{L}^{\psi}(u)=L(u)$. We then update the label $\widehat{L}^{\psi}(u)$ of every node $u$ as follows.
  • For every node $u$, the label $\ell(\rho(v))$ is added to $\widehat{L}^{\psi}(u)$, where
    $\ell(\rho(v))=\begin{cases}\rho(F)&\text{ if }\rho(F)\leq k\\ \infty&\text{ if }\rho(F)>k.\end{cases}$
  • If $u$ is a good node, then the label good is added $\widehat{L}^{\psi}(u)$.
@@
Let us suppose that there exists a is a lca-reduced $\widehat{\Lambda}^{\psi}$-formula $\alpha(\overline{x})$ that checks whether $\rho(F)>\rho(F,\overline{x})$ where $z$ is the ancestor at level $h$ of $x_{1}$. We can then define $\widehat{\psi}$ as the lca-reduced $\widehat{\Lambda}^{\psi}$-formula
and obtain that $\textsf{true}(\psi,F,L)=\textsf{true}(\widehat{\psi},F,\widehat{L}^{\psi})$ for every $\Lambda$-labeled forest $(F,L)$ of depth at most $d$. So, again, it is sufficient to show that $\alpha(\overline{x})$ exists.
@@
The fact that $\alpha(\overline{x})$ can be expressed as an lca-reduced $\widehat{\Lambda}^{\psi}$-formula follows from the same argument as for the formula $\alpha$ defined in Case 2, by modifying the definitions of $p$ and of the sets $X_{i}$. Indeed, observe that $F_{\xi}$ can contain at most $k$ roots. Therefore $\alpha$ is true if $\rho(F)>k$, which is equivalent the predicate $\textsf{lab}_{\infty}(x_{1})$. If $\rho(F)\geq k$ then the value of $\rho(F)$ is stored in the label of $x_{1}$. For every $p\leq k$, let $r_{1},\dots,r_{p}$ denote the roots of $F_{\xi}$. For every $i\in[p]$, let $X_{i}$ be the set of variables in $\overline{x}$ that are descendants of $r_{i}$. We obtain that checking whether $q=\rho(F)>\rho(F,\overline{x})$ is equivalent to check the existence of a set $S\in K(p,q-1)$ such that, for every $i\in[p]$, $X_{i}$ contains a good node if and only if $i\in S$, from which the expression of $\alpha$ follows. ∎
@@ 3.1.3 Quantifier Elimination of lca-Reduced Formulas in CONGEST
We now show that the quantifier elimination described in the proof of Lemma 4 can be computed in $\mathcal{O}(D)$ rounds in the CONGEST model.
Lemma 5.
Let $(G,L)$ be a $\Lambda$-labeled graph of diameter $D$, and let $F\in\mathcal{F}_{d}$ be a subgraph of $G$. Let $\varphi(\overline{x})=\exists\overline{y}\;\zeta(\overline{x},\overline{y})$ be a formula in $\text{FO}[\Lambda]$. There is a distributed algorithm performing in $O(D)$ round in the CONGEST model satisfying the following:
  • As input, every node $v\in V(G)$ is given $\Lambda$, $\varphi$, a set of labels $L(v)$, whether it belongs to $F$, whether it is a root of a tree in $F$, and which of its incident edges are in $F$;
  • As output, every node $v\in V(G)$ returns a set $\widehat{\Lambda}$, an lca-reduced formula $\widehat{\varphi}$, and its labels in a $\widehat{\Lambda}$-labeling $\widehat{L}$ such that $\textsf{true}(\varphi,F,L)=\textsf{true}(\widehat{\varphi},F,\widehat{L}).$
@@
Let us denote by $n$ the size of $G$. Let $T_{1},\dots,T_{p}$ by the trees in $F$, and let us denote by $r_{1},\dots,r_{p}$ the roots of these trees. The algorithm starts by performing a DFS traversal of each $T_{i}$ starting from $r_{i}$, in parallel for all $i=1,\dots,p$, whose objective is to compute the depth of each node in each tree. This takes $\mathcal{O}(d)$-rounds as $F\in\mathcal{F}_{d}$.
@@
Let $k=|\overline{x}|$. As for the proof of Lemma 4, we assume w.l.o.g. that $k\geq 1$, as otherwise we can create a dummy free variable. Also, we treat only the case where $|\overline{y}|=1$, as the general case $|\overline{y}|\geq 1$ can be obtained by iteratively eliminating the $|\overline{y}|$ quantifiers one by one. Let us define the set $I\subseteq~{}\textsf{Type}(k+1,D,\Lambda)$, and for each $\psi\in I$ the label sets $\widehat{\Lambda}^{\psi}$, lca-reduced formulas $\widehat{\psi}$ and $\hat{\Lambda}^{\psi}$-labeling $\widehat{L}^{\psi}$ of $F$ in the same way we defined them in Lemma 4. Furthermore, we define
@@
$\widehat{\Lambda}=\bigcup_{\psi\in I}\widehat{\Lambda}^{\psi},\;\widehat{\varphi}=\bigvee_{\psi\in I}\widehat{\psi},\text{ and, for every node $u\in F$, }\widehat{L}(u)=\bigcup_{\psi\in I}\widehat{L}^{\psi}(u).$
@@
We then extend $\widehat{L}$ to the nodes of $G\smallsetminus F$, by merely defining $\widehat{L}(v)=L(v)$ for all $v\notin V(F)$. Since $\textsf{Type}(k+1,d,\Lambda)$ is a finite set, it is enough to show that, for every $\psi\in I$, the label set $\widehat{\Lambda}^{\psi}$, and the formulas $\widehat{\psi}$ and $\widehat{L}^{\psi}$ can be computed in $\mathcal{O}(D)$ rounds. In fact, the definitions of the set $I$, $\widehat{\Lambda}$ and $\widehat{\varphi}$ do not depend in the instance $F$, so they can be computed at every node without any communication. It remains to compute, for any given $\psi\in I$, the value of $\widehat{L}^{\psi}(v)$ at each node $v$.
@@
Initially $\widehat{L}^{\psi}(v)=L(v)$ at every node. Let $\gamma$ and $\delta$ be the functions such that $\psi=\textsf{type}_{\gamma,\delta}$. Let $h_{x}$, $h_{y}$, and $h$ as defined in the proof of Lemma 4. Observe that every node knows $\varphi$, and thus every node can compute $\psi$, $h_{x}$, $h_{y}$, and $h$ without any communication. Furthermore, thanks to the DFS traversal performed in each $T_{i}$, every node knows whether its depth is $h$, $h_{x}$, $h_{y}$ or none of them.
We now revisit the three cases treated in the proof of Lemma 4.
Case 1: $\delta(1,k+1)\geq 0$ and $h_{y}=h$.
@@
In this case, the algorithm just needs to to identify the good nodes in $F$. To do so, for every $i\in[p]$, every node $v$ at depth $h_{y}$ in $T_{i}$ checks whether it is a candidate node. Every such nodes does not need any communication for this, as it knows its depth in $T_{i}$, as well as $L(v)$ and $\gamma(v)$. Then, the algorithm broadcasts ping messages (on a single bit) downward every tree, from every candidate node. These ping messages are forwarded by their descendants downward every $T_{i}$, for $h_{x}-h$ rounds. More precisely, each candidate node sends a ping as a single bit 1 to all its children, and, for $h-h_{x}-1$ rounds, each node that receives that ping from its parent forwards it to all its children. Finally, every node $u$ at depth $h_{x}$ that receives the ping know that their ancestor at level $h_{y}$ are candidates, and thus they mark themselves as good nodes. Moreover, every such node $u$ sets $\widehat{L}^{\psi}(u)=L(u)\cup\{\textsf{good}\}$. Every other node $u$ in $F$ sets $\widehat{L}^{\psi}(u)=L(u)$. All these operations consume $h_{x}-h=\mathcal{O}(d)$ rounds.
@@
This is the most demanding case. The algorithm consists in four phases. During the first phase, the algorithm aims at identifying the active roots. First, the candidate nodes are identified in the same way as in the previous case. By broadcasting a ping message upward each tree from the candidate nodes for $h_{y}$ rounds, each root becomes aware of the presence of a candidate node in its tree. The roots that receive a ping message mark themselves as active roots. This phase takes $\mathcal{O}(d)$ rounds. During the second phase, the algorithm aims at identifying the good children, by broadcasting a ping message downward from every active root. Every node $u$ that receives such ping message on this phase marks itself as good, and adds the label good to $\widehat{L}^{\psi}(u)$. This phase takes $\mathcal{O}(d)$ rounds. At this point, the algorithm constructs a DFS tree $T$ of $G$, rooted at node $r_{1}$. This procedure takes $\mathcal{O}(D)$ rounds. During the third phase, the root $r_{1}$ gets the number $\rho(F)$ of active roots in $F$, by counting upward the DFS tree $T$. More specifically, let us denote by $d_{T}=\mathcal{O}(D)$ the depth of $T$. The counting proceeds in $d_{T}+1$ round. For each node $u$ in $T$, let $a(u)$ be $1$ if $u$ is an active root, and 00 otherwise. In the first round, each node $u$ at depth $d_{T}$ set $\textsf{val}(u)=a(u)$ and sends $\textsf{val}(u)$ to its parent in $T$. Them for each $i\in[d_{t}]$, at round $i+1$, every node $u$ of $T$ at depth $d_{T}-i$ computes $\textsf{val}(u)=a(u)+\sum_{v\in C(u)}\textsf{val}(u)$, where $C(u)$ is the set of children of $u$ in $T$, and sends $\textsf{val}(u)$ to its parent. At the end of this phase, $\textsf{val}(r)=\rho(F)$. Observe that, for every node $u$, $\textsf{val}(u)\leq n$, and thus this value can be communicated in one round under the CONGEST model. Therefore, the third phase runs in $\mathcal{O}(D)+d_{T}+1=\mathcal{O}(D)$ rounds. During the fourth phase, the root $r_{1}$ broadcast $\rho(F)$ downward the tree $T$, to every node in $G$. Then, each node $u$ in $F$ adds the label $\ell(\rho(F))$ to $\widehat{F}^{\psi}(u)$. This phase takes $d_{T}=O(D)$ rounds.
We conclude that for every $\psi\in I$, every node $u$ can compute $L^{\psi}(u)$, in $\mathcal{O}(D)$ rounds. Therefore, the total number of rounds sufficient for computing $\widehat{L}$ is $|I|\cdot D\leq c\cdot 2^{k|\Lambda|}(d+2)^{k^{2}}\cdot D$, where $c$ is a constant not depending on $d$, $\Lambda$ or $\psi$ or $D$. ∎
3.2 Graphs of Bounded Treedepth

### v2

@@ 3.1.2 Quantifier Elimination of lca-Expressed Formulas
$\alpha(\overline{x})=\bigvee_{q\in[p]}\bigg{[}\bigg{(}\bigvee_{q\leq m\leq k+1}\textsf{lab}_{m}(x_{s})\bigg{)}\wedge\bigvee_{S\in K(p,q-1)}\bigg{(}\bigwedge_{i\in S}\bigvee_{j\in X_{i}}\textsf{lab}_{\textsf{good}}(x_{j})\wedge\bigwedge_{i\notin S}\bigwedge_{j\in X_{i}}\neg\textsf{lab}_{\textsf{good}}(x_{j})\bigg{)}\bigg{]}.$
We claim that $\textsf{true}(\psi,F,\ell)=\textsf{true}(\widehat{\psi},F,\widehat{\ell}^{\psi})$ for every $\Lambda$-labeled forest $(F,\ell)\in\mathcal{F}_{d}$. Indeed, observe first that $\kappa(z,\overline{x})\leq k$ as the forest $F_{\xi}$ has degree at most $k$ (because it has at most $k$ leaves). Therefore, $\alpha(\overline{x})$ holds whenever $\kappa(z)>k$, which is equivalent to the predicate $\textsf{lab}_{k+1}(x_{s})$. If $\kappa(z)\leq k$ then the value of $\kappa(z)$ is stored in the label of $x_{s}$. Let us denote by $z_{1},\dots,z_{p}$ the $p\leq k$ children of $z$ in $F_{\psi}$. For every $i\in[p]$, let us denote by $X_{i}$ the set of variables in $\overline{x}$ that are descendants of $z_{i}$. For every $q\in[p]$, we also denote by $K(p,q)$ the set of all subsets of $[p]$ of at most $q$ elements. If $q=\kappa(z)$ and $t=\kappa(z,\overline{x})$, then checking $q>t$ can be done by checking the existence of set $S\in K(p,q-1)$ satisfying that, for every $i\in S$, $X_{i}$ contains a good node, and, for every $i\notin S$, $X_{i}$ does not contain a good node.
Case 3: $\delta(s,k+1)=-1$.
In this case, the node $y$ in $F_{\psi}$ has no common ancestors with any of the nodes $x_{1},\dots,x_{k}$. In other words, $y$ is in a tree of $F_{\psi}$ that is different from any of the trees containing $x_{1},\dots,x_{k}$. Let $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$. Like in previous cases, a node $u$ at depth $h_{y}$ is called a candidate if $\ell(u)=\gamma(k+1)$. A root $r$ of $F$ is said to be active if some descendant of $r$ is a candidate node. A node is called good if it is a descendant of an active root.
@@
Let $\overline{v}=(v_{1},\dots,v_{k})$ be a set of vertices of $F$, and let $\rho(F)$ be the number of active roots in $F$. Similarly, let $\rho(F,\overline{v})$ be the number of active roots in $F[\overline{v}]$. It follows that $(F,\ell)\models\exists y,\psi(\overline{v},y)$ if and only if $F[\overline{v}]$ is isomorphic to $F_{\xi}$, $\ell(v_{i})=\gamma(i)$ for each $i\in[k]$, and $F$ contains an active root not contained in $F[\overline{v}]$. The first two conditions can be checked using $\xi$, and the latter is equivalent to $\rho(F)>\rho(F,\overline{v})$.
@@
Let $\widehat{\Lambda}=\Lambda\cup[0,k+1]\cup\{\textsf{good}\}$ be a set of labels, and let us consider the $\widehat{\Lambda}^{\psi}$-labeling ^$\widehat{\ell}^{\psi}$ of $F$ defined as follows. Initially, for every node $u$ in $F$, let us set ^$\widehat{\ell}^{\psi}(u)=\ell(u)$. We then update the label ^$\widehat{\ell}^{\psi}(u)$ of every node $u$ as follows.
  • For every node $u$, the label $\lambda$ is added to ^$\widehat{\ell}^{\psi}(u)$, where
    $\lambda=\begin{cases}\rho(F)&\text{ if }\rho(F)\leq k\\ k+1&\text{ if }\rho(F)>k.\end{cases}$
  • If $u$ is a good node, then the label good is added ^$\widehat{\ell}^{\psi}(u)$.
@@
Let us suppose that there exists an lca-reduced $\widehat{\Lambda}^{\psi}$-formula $\alpha(\overline{x})$ that checks whether $\rho(F)>\rho(F,\overline{x})$ where $z$ is the ancestor at level $h$ of $x_{1}$. We can then define $\widehat{\psi}$ as the lca-reduced $\widehat{\Lambda}^{\psi}$-formula
$\widehat{\psi}(\overline{x})=\textsf{type}_{\widehat{\gamma},\widehat{\delta}}(\overline{x})~{}\wedge~{}\xi(\overline{x}),$
where
$\xi(\overline{x})=\bigvee_{q\in[p]}\bigg{[}\bigg{(}\bigvee_{q\leq m\leq k+1}\textsf{lab}_{m}(x_{1})\bigg{)}\wedge\bigvee_{S\in K(p,q-1)}\bigg{(}\bigwedge_{i\in S}\bigvee_{j\in X_{i}}\textsf{lab}_{\textsf{good}}(x_{j})\wedge\bigwedge_{i\notin S}\bigwedge_{j\in X_{i}}\neg\textsf{lab}_{\textsf{good}}(x_{j})\bigg{)}\bigg{]}$ (1)
@@
We claim that $\textsf{true}(\psi,F,\ell)=\textsf{true}(\widehat{\psi},F,\widehat{\ell}^{\psi})$ for every $\Lambda$-labeled forest $(F,\ell)\in\mathcal{F}_{d}$. Indeed, observe that $F_{\xi}$ can contain at most $k$ roots. Therefore $\alpha$ is true if $\rho(F)>k$, which is equivalent the predicate $\textsf{lab}_{k+1}(x_{1})$. If $\rho(F)\leq k$ then the value of $\rho(F)$ is stored in the label of $x_{1}$. Let $r_{1},\dots,r_{p}$ denote the $p\leq k$ roots of $F_{\xi}$. For every $i\in[p]$, let $X_{i}$ be the set of variables in $\overline{x}$ that are descendants of $r_{i}$. We obtain that checking whether $q=\rho(F)>\rho(F,\overline{x})$ is equivalent to check the existence of a set $S\in K(p,q-1)$ such that, for every $i\in[p]$, $X_{i}$ contains a good node if and only if $i\in S$, from which the expression of $\xi$ follows. ∎
Finally, the following lemma is a refined version of Lemma 3.7 enforcing locality.
Lemma 3.8.
Let $d\in\mathbb{N}$, $r\geq d$, and $\varphi(\overline{x})=\exists\overline{y}\,\big{(}\zeta(\overline{x},\overline{y})\wedge\mathrm{ball}(\{\overline{x},\overline{y}\},x_{1},r)\big{)}$ be an $r$-local lca-expressed existential formula in $\mathrm{FO}[\Lambda]$, i.e., $\zeta$ is a lca-reduced formula in $\mathrm{FO}[\Lambda]$. It holds that $\varphi$ is $d$-local reducible on $\mathcal{F}_{d}[\Lambda]$.
Proof.
Since all trees in the forests $F\in\mathcal{F}_{d}$ have diameter at most $d$, and since $r\geq d$, we get that $\mathrm{ball}(z,x_{1},r)$ holds if and only if $x_{1}$ and $z$ are in the same tree, which is equivalent to $\neg\textsf{lca}_{-1}(x_{1},z)$. We define:
$\varphi_{1}(\overline{x})=\exists\overline{y},\zeta(\overline{x},\overline{y})\wedge\bigwedge_{z\in\overline{x},\overline{y}}\neg\textsf{lca}_{-1}(x_{1},z)$
We have $\textsf{true}(\varphi,F,\ell)=\textsf{true}(\varphi_{1},F,\ell)$. We can now apply Lemma 3.7 on $\varphi_{1}$. (Note that there were three cases the proof of Lemma 3.7, but the third case cannot happen here as this case corresponds to the case where two variables are not in the same connected components of $F$, which is forbidden by the terms of the form $\neg\textsf{lca}_{-1}(x_{1},z)$.) By Lemma 3.7, we obtain an lca-reduced formula $\widehat{\varphi}$ and a labeling $\widehat{\ell}$ such that $\textsf{true}(\varphi_{1},F,\ell)=\textsf{true}(\widehat{\varphi},F,\widehat{\ell})$. Now, as mentioned above, each $x_{i}$ is in the same component as $x_{1}$. Since each tree has diameter at most $d$, this mere fact implies that $\mathrm{ball}(x_{i},x_{1},d)$ holds for every $i$. As a consequence, $\textsf{true}(\widehat{\varphi},F,\widehat{\ell})=\textsf{true}(\widehat{\varphi}\wedge\mathrm{ball}(\overline{x},x_{1},d),F,\widehat{\ell})$. ∎
@@ 3.1.3 Quantifier Elimination of lca-Reduced Formulas in CONGEST
We now show that the quantifier elimination on local formulas described in the proof of Lemma 3.8 can be computed in $\mathcal{O}(1)$ rounds in the CONGEST model.
Lemma 3.9.
Let $\Lambda$ be a finite set of labels, $d\in\mathbb{N}$ and $r\geq d$. For every $r$-local lca-expressed existential formula $\varphi(\overline{x})=\exists\overline{y}\;\big{(}\zeta(\overline{x},\overline{y}\wedge\mathrm{ball}(\{\overline{x},\overline{y}\},x_{1},r))\big{)}$ in $\mathrm{FO}[\Lambda]$, there exists a set of labels $\widehat{\Lambda}$, a lca-reduced formula $\widehat{\varphi}$ and a distributed algorithm performing in $\mathcal{O}(1)$ round in the CONGEST model satisfying the following. For every $\Lambda$-labeled graph $(G,\ell)$, every subgraph $F\in\mathcal{F}_{d}$ of $G$:
  • every node $v\in V(G)$ is given as input: a set of labels $\ell(v)\subseteq\Lambda$, whether it belongs to $F$, and, if applicable, its depth and parent in $F$;
  • every node $v\in V(G)$ outputs a set of labels ^$\widehat{\ell}(v)\subseteq\widehat{\Lambda}$ such that $\widehat{\Lambda},\widehat{\varphi},\widehat{\ell}$ form a $d$-local reduction of $\Lambda,\varphi,\ell$.
@@
Let $k=|\overline{x}|$. As for the proof of Lemma 3.7, we assume w.l.o.g. that $k\geq 1$, and we treat only the case where $|\overline{y}|=1$ (as the general case $|\overline{y}|\geq 1$ can be obtained by iteratively eliminating the $|\overline{y}|$ quantifiers one by one). In the same way they are defined in Lemma 3.7, let us define the set $I\subseteq~{}\textsf{Type}(k+1,D,\Lambda)$, and, for each $\psi\in I$, the label sets $\widehat{\Lambda}^{\psi}$, the lca-reduced formula $\widehat{\psi}$, and the $\hat{\Lambda}^{\psi}$-labeling ^$\widehat{\ell}^{\psi}$ of $F$. Furthermore, we define
@@
$\widehat{\Lambda}=\bigcup_{\psi\in I}\widehat{\Lambda}^{\psi},\;\widehat{\varphi}=\bigvee_{\psi\in I}\widehat{\psi},\text{ and, for every node $u\in F$, }\widehat{\ell}(u)=\bigcup_{\psi\in I}\widehat{\ell}^{\psi}(u).$
@@
We then extend $\widehat{\ell}$ to the nodes of $G\smallsetminus F$ by merely defining ^$\widehat{\ell}(v)=\ell(v)$ for all $v\notin V(F)$. Since $\textsf{Type}(k+1,d,\Lambda)$ is a finite set, it is enough to show that, for every $\psi\in I$, the label set $\widehat{\Lambda}^{\psi}$, and the formulas $\widehat{\psi}$ and ^$\widehat{\ell}^{\psi}$ can be computed in $\mathcal{O}(1)$ rounds. In fact, the definitions of the set $I$, $\widehat{\Lambda}$, and $\widehat{\varphi}$ do not depend in the sub-forest $F$ of $G$, so $I$, $\widehat{\Lambda}$, and $\widehat{\varphi}$ can be computed at every node without any communication. It remains to show how to compute, for any given $\psi\in I$, the value of ^$\widehat{\ell}^{\psi}(v)$ at all nodes $v$ in a constant number of rounds.
@@
Initially the algorithm sets ^$\widehat{\ell}^{\psi}(v)=\ell(v)$ at every node. Let $\gamma$ and $\delta$ be the functions such that $\psi=\textsf{type}_{\gamma,\delta}$. Let $s$, $h_{s}$, $h_{y}$, and $h$ be as defined in the proof of Lemma 3.7. Observe that every node knows $\varphi$, and thus every node can compute $\psi$, $h_{s}$, $h_{y}$, and $h$ without any communication. Furthermore, every node knows whether its depth is equal to $h$, $h_{s}$, or $h_{y}$, or none of these three values.
We now revisit the three cases treated in the proof of Lemma 3.7.
Case 1: $\delta(s,k+1)\geq 0$ and $h_{y}=h$.
@@
In this case, the algorithm just needs to identify the good nodes in $F$. To do so, for every $i\in[p]$, every node $v$ at depth $h_{y}$ in $T_{i}$ checks whether it is a candidate node. Every such nodes does not need any communication for this, as it knows its depth in $T_{i}$, as well as $\ell(v)$ and $\gamma(v)$. Then, the algorithm broadcasts ping messages (on a single bit) downward every tree, from every candidate node. These ping messages are forwarded by their descendants downward every $T_{i}$, for $h_{s}-h$ rounds. More precisely, each candidate node sends a ping as a single bit 1 to all its children, and, for $h-h_{s}-1$ rounds, each node that receives that ping from its parent forwards it to all its children. Finally, every node $u$ at depth $h_{s}$ that receives the ping know that their ancestor at level $h_{y}$ are candidates, and thus they mark themselves as good nodes. Moreover, every such node $u$ sets ^$\widehat{\ell}^{\psi}(u)=\ell(u)\cup\{\textsf{good}\}$. Every other node $u$ in $F$ sets ^$\widehat{\ell}^{\psi}(u)=\ell(u)$. All these operations consume $h_{s}-h=\mathcal{O}(d)$ rounds.
@@
As mentioned in the proof of Lemma 3.8, this case cannot happen for local formulas.
We conclude that for every $\psi\in I$, every node $u$ can compute $\ell^{\psi}(u)$, in $\mathcal{O}(d)$ rounds. Therefore, the total number of rounds sufficient for computing $\widehat{\ell}$ is $\mathcal{O}(d\cdot|I|)=\mathcal{O}(d\cdot 2^{k|\Lambda|}(d+1)^{k^{2}})$. ∎
3.2 Quantifier Elimination on Graphs of Bounded Treedepth